OpenStudy (anonymous):

Write the equation 2x - 3y = 6 in slope-intercept form.

OpenStudy (anonymous):

Hi welcome to Openstudy! slope intercept form is y=mx + b, where m is the slope and be is the y-intercept. so we have to evaluate the given equation in terms of y.

OpenStudy (anonymous):

first thing you have to do is to subtract both sides by 2x. can you do that for me?

OpenStudy (anonymous):

3y=6-2x

OpenStudy (anonymous):

right :) now divide both sides by 3

OpenStudy (anonymous):

y=2-2x?

OpenStudy (anonymous):

almost right.. you forgot to divide 2x by 3

OpenStudy (anonymous):

so the equation will be \(y=2-\frac{2x}{3}\) let's rearrange it in terms of y=mx+b the final equation will be \(y=-\frac{2x}{3 }+2\)

OpenStudy (anonymous):

any confusions??

OpenStudy (anonymous):

Well the answer choice says that exactly except its minus 2 at the end? or the other option is plus 3 at the end

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

oh yeah, we should divide both sides by negative 3 and not positive 3. sorry :/